How may I solve these systems by adding or subtracting:
2x-11y=15
5x+3y=7

Answer by scott8148(6628)

multiplying 1st equation by 3 __ 6x-33y=45

multiplying 2nd equation by 11 __ 55x+33y=77

ADDING the new equations __ 61x=122 __ x=2

substituting __ 5(2)+3y=7 __ 10+3y=7 __ 3y=-3 __ y=-1